# Streaming Operator Inference for Model Reduction of

### v1

@@ 1 Introduction
The ability to use data to learn accurate and efficient reduced models for complex dynamical systems is crucial across scientific and engineering applications, as these models facilitate rapid simulations for many-query computations such as uncertainty quantification, inverse problems, optimization, and control.
Operator Inference (OpInf) Peherstorfer and Willcox [2016], Kramer et al. [2024] refers to a family of methods that learn such efficient reduced models from data by fitting reduced model operators to minimize the residual of the reduced system equations. OpInf assumes a polynomial reduced model, where the learned operators are coefficient matrices representing polynomial nonlinearities. The formulation hinges on two key steps: (i) dimension reduction by projecting data onto a reduced basis constructed via Proper Orthogonal Decomposition (POD), which computes an optimal low-dimensional approximation via Singular Value Decomposition (SVD), and (ii) solving a linear least-squares (LS) problem to compute the reduced operators by fitting them to all projected data at once. The first step emanates from classical projection-based model reduction, which projects the full-dimensional model onto a low-dimensional subspace to yield accurate yet computationally efficient reduced models Benner et al. [2015], while the second step leverages data-driven learning to infer reduced operators without intrusive access to the underlying code or full model operators, and thereby making OpInf a non-intrusive method.
@@
Another related work, Streaming weak-SINDy Russo et al. [2025], learns weak-form partial differential equation (PDE) solutions in a streaming fashion. The method accumulates feature and target matrices via streaming integration with streaming POD, deferring regression to an offline stage where sequential thresholding LS requires solving the batch LS problem repeatedly until sparsity is achieved, whereas we consider online updating of the LS solution via RLS along with different combinations of algorithms for streaming reduced operator learning. All existing streaming methods introduced so far show growing recognition that incremental learning is essential for scalability and real-time adaptation. However, we emphasize that no existing approach combines dimension reduction via incremental basis updates with direct learning of reduced polynomial operators from streaming data, which is the contribution of our Streaming OpInf framework.
@@ 2.1 The Operator Inference Problem
    where the symmetric positive definite matrix $\boldsymbol{\Gamma}\in\mathbb{R}^{d\times d}$ with $d=r+r^{2}+m+1$ defines a Tikhonov regularization term McQuarrie et al. [2021].
@@
If $\mathbf{D}$ has full column rank, this overdetermined system has unique solution $\mathbf{O}=\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{R}}\mskip-1.5mu}\mskip 1.5mu$.
Standard OpInf is a batch method built on standard direct computation of the SVD and direct solution of the LS problem. These approaches load and store the entire data set in working memory, which is infeasible for real-time streaming data settings and large-scale problems. While distributed computing Farcaş et al. [2025] and domain decomposition Farcas et al. [2024] address the memory bottleneck, we pursue an alternative by adapting OpInf to the streaming setting, where snapshots $\mathbf{x}_{1},\ldots,\mathbf{x}_{k}$ are processed incrementally as they arrive. This requires two key ingredients: (i) incremental basis construction via iSVD, and (ii) recursive operator learning via RLS which we now introduce.
@@ 2.2 Streaming SVD Methods
We seek to address the high memory cost of $O(n(K+r))$ in standard SVD for storing the data and basis, by introducing two streaming SVD algorithms, Baker’s iSVD Baker et al. [2012] and SketchySVD Tropp et al. [2019a].
@@ 2.2.1 Baker’s iSVD Algorithm
In this work, we employ Baker’s iSVD framework Baker et al. [2012], which builds on these identities (8) and (9) while addressing two key challenges in standard iSVD methods: determining whether new snapshots warrant basis updates (typically via residual norm $p$ thresholding Brand [2002]) and maintaining numerical stability against accumulated errors (often requiring ad-hoc reorthogonalization Oxberry et al. [2017], Brand [2002]). Baker’s algorithm eliminates update determination by enforcing strict orthonormality at every iteration via double orthogonalization: after projecting the new data vector, it applies Gram-Schmidt orthogonalization followed by QR decomposition on the residual, $\mathbf{x}_{\perp}p\leftarrow\mathsf{qr}(\mathbf{x}_{\perp})$. This approach, similar to Zha and Simon [1999], ensures numerical accuracy without additional reorthogonalization steps. For completeness, we summarize Baker’s iSVD algorithm in Algorithm 1.
@@ 2.2.2 Randomized SketchySVD Algorithm
Let the sketch sizes be denoted $q$ and $s$, which control the dimensions of the compressed representations and determine the accuracy-memory tradeoff. SketchySVD maintains three randomized sketches $\mathcal{X}_{\mathsf{range}}\in\mathbb{R}^{n\times q},\,\mathcal{X}_{\mathsf{corange}}\in\mathbb{R}^{q\times K},\,\mathcal{X}_{\mathsf{core}}\in\mathbb{R}^{s\times s}$ that capture the range, corange, and core information (information to improve the estimate of the singular values and vectors) of the data matrix $\mathbf{X}$, respectively Tropp et al. [2019a]. These sketches are formed as products of $\mathbf{X}$ with four fixed random reduction maps $\boldsymbol{\Upsilon}\in\mathbb{R}^{q\times n},\,\boldsymbol{\Omega}\in\mathbb{R}^{q\times K},\,\boldsymbol{\Xi}\in\mathbb{R}^{s\times n},\,\boldsymbol{\Psi}\in\mathbb{R}^{s\times K}$:
@@
After processing all snapshots, the algorithm computes QR decompositions of the range and corange sketches, forms a small core matrix $\mathbf{C}\in\mathbb{R}^{s\times s}$, computes its SVD, and combines these results to construct the approximate SVD of $\mathbf{X}$. To ensure a truncated SVD with sufficiently tight theoretical error bounds, the recommended sketch sizes are $q=4r+1$ and $s=2q+1$ Tropp et al. [2019a]. The SketchySVD algorithm is summarized in Algorithm 2, with comprehensive details available in Tropp et al. [2019a].
@@ 2.2.3 Cost and Error Analysis
To assess suitability for our proposed method, we analyze memory and computational costs as well as sources of errors of Baker’s iSVD and SketchySVD, comparing them to standard batch SVD in Table 1.
@@ 2.3 Recursive Least-Squares Methods
We introduce two RLS algorithms: standard RLS and a more numerically stable incremental QR-decomposition Recursive Least Squares (iQRRLS) in Sections 2.3 and 2.3.2.
@@ 2.3.2 Inverse QR-Decomposition RLS Algorithm
While the standard RLS algorithm 3 is effective for recursive parameter estimation, it suffers from numerical instabilities when updating the inverse correlation matrix $\mathbf{P}_{k}$ in finite-precision arithmetic. These instabilities arise because updating $\mathbf{P}_{k}$ involves subtracting two positive definite matrices, which can destroy positive definiteness and cause catastrophic cancellation when nearly equal large values are subtracted Sayed [2011]. To address these numerical instabilities, QR-decomposition RLS algorithms organize iterated quantities and incoming data into arrays that are factorized using QR decomposition to update the necessary variables in a numerically stable manner.
These algorithms rely on transformations of the form $\mathcal{A}_{k}\boldsymbol{\Theta}_{k}=\mathcal{B}_{k}$, where $\mathcal{A}_{k}$ is the pre-array containing previous quantities and new data, $\boldsymbol{\Theta}_{k}$ is a unitary matrix (e.g., a product of Givens rotations or Householder reflections), and $\mathcal{B}_{k}$ is the triangular post-array containing updated quantities. This transformation is essentially a QR decomposition of $\mathcal{A}_{k}^{\top}$, leading to two main variants: iQRRLS and QR-decomposition Recursive Least Squares (QRRLS), which differ in whether they update the inverse correlation matrix $\mathbf{P}_{k}$ or the correlation matrix $\sum_{i=1}^{k}\mathbf{d}_{i}^{\top}\mathbf{d}_{i}$ Sayed [2011], Haykin [2014].
The iQRRLS algorithm specifically addresses numerical instability by propagating the square root of the inverse correlation matrix $\mathbf{P}_{k}^{1/2}$ rather than $\mathbf{P}_{k}$ itself to ensure positive definiteness, using the Cholesky decomposition $\mathbf{P}_{k}=\mathbf{P}_{k}^{1/2}(\mathbf{P}_{k}^{1/2})^{\top\kern-0.5pt}$ where $\mathbf{P}_{k}^{1/2}$ is lower triangular. iQRRLS then organizes the transposed square root matrix from the previous iteration, denoted with shorthand notation $\mathbf{P}_{k-1}^{\top/2}=(\mathbf{P}_{k-1}^{1/2})^{\top\kern-0.5pt}$, and incoming data into the pre-array
@@
and then performs QR decomposition to obtain the post-array from which updated quantities can be extracted:
$c_{k}^{-1/2}\leftarrow\mathcal{B}_{k}^{\top}(1,1),\quad\mathbf{g}_{k}c_{k}^{-1/2}\leftarrow\mathcal{B}_{k}^{\top}(1,2:d+1),\quad\mathbf{P}_{k}^{1/2}\leftarrow{(\mathcal{B}_{k}^{\top}(2:d+1,2:d+1))}^{\top},$
where $(i,j)$ denotes the matrix entry at row $i$ and column $j$. We summarize iQRRLS in Algorithm 4.
@@ 3.1 Streaming Reformulation of the Learning Step
To circumvent this additional data pass and its associated cost, we introduce a reformulation that expresses the LS data matrices directly in terms of the truncated SVD matrices obtained from iSVD, thereby avoiding explicit projection entirely. This reformulation approach yields the iSVD-LS and iSVD-RLS paradigms.
@@
$\begin{gathered}\mathbf{X}_{k}\odot\mathbf{X}_{k}\approx(\mathbf{V}_{r_{k}}\otimes\mathbf{V}_{r_{k}})(\boldsymbol{\Sigma}_{r_{k}}\otimes\boldsymbol{\Sigma}_{r_{k}})(\mathbf{W}_{r_{k}}\mathchoice{\mathbin{\vtop{\halign{#\cr$\displaystyle=$\crcr$\displaystyle\ovoid$\crcr}}}}{\mathbin{\vtop{\halign{#\cr$\textstyle=$\crcr$\textstyle\ovoid$\crcr}}}}{\mathbin{\vtop{\halign{#\cr$\scriptstyle=$\crcr$\scriptstyle\ovoid$\crcr}}}}{\mathbin{\vtop{\halign{#\cr$\scriptscriptstyle=$\crcr$\scriptscriptstyle\ovoid$\crcr}}}}\mathbf{W}_{r_{k}})^{\top\kern-0.5pt}\\ \implies\hat{\mathbf{X}}_{k}\odot\hat{\mathbf{X}}_{k}=(\boldsymbol{\Sigma}_{r_{k}}\otimes\boldsymbol{\Sigma}_{r_{k}})(\mathbf{W}_{r_{k}}\mathchoice{\mathbin{\vtop{\halign{#\cr$\displaystyle=$\crcr$\displaystyle\ovoid$\crcr}}}}{\mathbin{\vtop{\halign{#\cr$\textstyle=$\crcr$\textstyle\ovoid$\crcr}}}}{\mathbin{\vtop{\halign{#\cr$\scriptstyle=$\crcr$\scriptstyle\ovoid$\crcr}}}}{\mathbin{\vtop{\halign{#\cr$\scriptscriptstyle=$\crcr$\scriptscriptstyle\ovoid$\crcr}}}}\mathbf{W}_{r_{k}})^{\top\kern-0.5pt}=\boldsymbol{\Sigma}_{r_{k}}\mathbf{W}_{r_{k}}^{\top}\odot\boldsymbol{\Sigma}_{r_{k}}\mathbf{W}_{r_{k}}^{\top},\end{gathered}$ (14)
[…]
Using (13) and (14), the data matrix $\mathbf{D}_{k}$ from (5) can be reformulated solely in terms of the rank-$r$ truncated matrices $\boldsymbol{\Sigma}_{r}=\boldsymbol{\Sigma}_{r_{k}}(1:r,1:r)$ and $\mathbf{W}_{r}=\mathbf{W}_{r_{k}}(:,1:r)$:
$\mathbf{D}_{k}=\begin{bmatrix}\mathbf{W}_{r}\boldsymbol{\Sigma}_{r}&\mathbf{W}_{r}\boldsymbol{\Sigma}_{r}\mathchoice{\mathbin{\vtop{\halign{#\cr$\displaystyle=$\crcr$\displaystyle\ovoid$\crcr}}}}{\mathbin{\vtop{\halign{#\cr$\textstyle=$\crcr$\textstyle\ovoid$\crcr}}}}{\mathbin{\vtop{\halign{#\cr$\scriptstyle=$\crcr$\scriptstyle\ovoid$\crcr}}}}{\mathbin{\vtop{\halign{#\cr$\scriptscriptstyle=$\crcr$\scriptscriptstyle\ovoid$\crcr}}}}\mathbf{W}_{r}\boldsymbol{\Sigma}_{r}&\mathbf{U}_{k}^{\top}&\mathbf{1}_{k}\end{bmatrix}$,
noting that $\mathbf{W}_{r}\boldsymbol{\Sigma}_{r}\mathchoice{\mathbin{\vtop{\halign{#\cr$\displaystyle=$\crcr$\displaystyle\ovoid$\crcr}}}}{\mathbin{\vtop{\halign{#\cr$\textstyle=$\crcr$\textstyle\ovoid$\crcr}}}}{\mathbin{\vtop{\halign{#\cr$\scriptstyle=$\crcr$\scriptstyle\ovoid$\crcr}}}}{\mathbin{\vtop{\halign{#\cr$\scriptscriptstyle=$\crcr$\scriptscriptstyle\ovoid$\crcr}}}}\mathbf{W}_{r}\boldsymbol{\Sigma}_{r}=\boldsymbol{\Sigma}_{r}\mathbf{W}_{r}^{\top}\odot\boldsymbol{\Sigma}_{r}\mathbf{W}_{r}^{\top}$ since the Khatri-Rao product equals the transposed face-splitting product.
@@ 3.3.1 Errors from Streaming SVD
Baker’s iSVD produces deterministic errors that accumulate with each truncation Fareed and Singler [2020], Baker et al. [2012], while SketchySVD produces probabilistic errors controlled by sketch parameters Tropp et al. [2019a]. Both methods exhibit increased errors when spectral gaps diminish, as discussed in Section 2.2.3. These errors propagate to the learned operators through the LS data matrices. Specifically, the iSVD-LS/RLS paradigms construct data matrices (15) using right singular vectors $\mathbf{W}_{r}$ and singular values $\boldsymbol{\Sigma}_{r}$, which carry larger errors than left singular vectors according to (10). In contrast, the iSVD-Project-LS/RLS paradigms project data onto the POD basis using left singular vectors $\mathbf{V}_{r}$, yielding lower operator error at additional computational cost. We formalize this error propagation in the following lemmas and theorem.
The first lemma establishes a general perturbation bound for the learned operators in terms of perturbations to the data matrices.
@@
Let $\widetilde{\mathbf{D}}=\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu+[\mathbf{E}_{D}^{\top},\mathbf{0}_{d\times d}]^{\top\kern-0.5pt}$ and $\widetilde{\mathbf{R}}=\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{R}}\mskip-1.5mu}\mskip 1.5mu+[\mathbf{E}_{R}^{\top},\mathbf{0}_{r\times d}]^{\top\kern-0.5pt}$ denote perturbed versions, where the perturbations $\mathbf{E}_{D}\in\mathbb{R}^{K\times d}$ and $\mathbf{E}_{R}\in\mathbb{R}^{K\times r}$ arise from approximate SVD factors. Then the operator error between the batch solution $\mathbf{O}=\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{R}}\mskip-1.5mu}\mskip 1.5mu$ and the streaming solution $\widetilde{\mathbf{O}}=\widetilde{\mathbf{D}}^{\dagger}\widetilde{\mathbf{R}}$ satisfies
@@
The second lemma bounds the data matrix perturbations $\mathbf{E}_{D}$ and $\mathbf{E}_{R}$ in terms of the singular vector and singular value errors from streaming SVD. Let $\{\mathbf{v}_{j},\sigma_{j},\mathbf{w}_{j}\}_{j=1}^{r}$ and $\{\widetilde{\mathbf{v}}_{j},\widetilde{\sigma}_{j},\widetilde{\mathbf{w}}_{j}\}_{j=1}^{r}$ denote the singular triplets from exact and approximate rank-$r$ SVD, respectively, where the singular vectors are orthonormal.
Lemma 3.2 (Data matrix perturbation bounds).
@@
Let $\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu$ and $\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{R}}\mskip-1.5mu}\mskip 1.5mu$ be constructed using the exact rank-$r$ SVD of $\mathbf{X}$, and let $\widetilde{\mathbf{D}}$ and $\widetilde{\mathbf{R}}$ be constructed using approximate factors $\widetilde{\mathbf{V}}_{r}$, $\widetilde{\boldsymbol{\Sigma}}_{r}$, $\widetilde{\mathbf{W}}_{r}$ from streaming SVD. Assume the time-derivative data is computed via finite differences, i.e., $\dot{\mathbf{X}}=\mathbf{X}\boldsymbol{\Delta}_{K}$. Suppose $\|\mathbf{X}-\widetilde{\mathbf{X}}\|_{2}\leq\varepsilon$, which implies $|\sigma_{j}-\widetilde{\sigma}_{j}|\leq\varepsilon$ for all $j=1,\ldots,r$. Let $\|\mathbf{v}_{j}-\widetilde{\mathbf{v}}_{j}\|_{2}\leq\varepsilon_{v,j}$ and $\|\mathbf{w}_{j}-\widetilde{\mathbf{w}}_{j}\|_{2}\leq\varepsilon_{w,j}$ for all $j$. Define the cumulative singular vector errors
$\tau_{v}=\Bigl(\sum_{j=1}^{r}\varepsilon_{v,j}^{2}\Bigr)^{1/2},\qquad\tau_{w}=\Bigl(\sum_{j=1}^{r}\varepsilon_{w,j}^{2}\Bigr)^{1/2}.$ (17)
Then for the iSVD-Project-LS/RLS paradigms (projection onto POD basis),

### v2

@@ 1 Introduction
The ability to use data to learn accurate and efficient reduced models for complex dynamical systems is crucial across scientific and engineering applications, as these models facilitate rapid simulations for many-query computations such as uncertainty quantification, inverse problems, optimization, and control. Operator Inference (OpInf) Peherstorfer and Willcox [2016], Kramer et al. [2024] refers to a family of methods that learn such efficient reduced models from data by fitting reduced model operators to minimize the residual of the reduced system equations. Assuming a polynomial reduced model, OpInf learns the coefficient matrices representing polynomial nonlinearities. OpInf computes such reduced operators following two key steps: (i) dimension reduction by projecting data onto a reduced basis constructed via Proper Orthogonal Decomposition (POD), which computes an optimal low-dimensional approximation via Singular Value Decomposition (SVD), and (ii) solving a linear least-squares (LS) problem to compute the reduced operators by fitting them to all projected data at once. The first step emanates from classical projection-based model reduction, which projects the full-dimensional model onto a low-dimensional subspace to yield accurate yet computationally efficient reduced models Benner et al. [2015], while the second step leverages data-driven learning to infer reduced operators without intrusive access to the underlying code or full model operators, and thereby making OpInf a non-intrusive method.
@@
Another related work, Streaming weak-SINDy Russo et al. [2025], learns weak-form partial differential equation (PDE) solutions. The method accumulates feature and target matrices via streaming integration with streaming POD, deferring regression to an offline stage where sequential thresholding LS requires solving the batch LS problem repeatedly until sparsity is achieved, whereas we consider online updating of the LS solution via RLS along with different combinations of algorithms for streaming reduced operator learning. All existing streaming methods introduced so far show growing recognition that incremental learning is essential for scalability and real-time adaptation. However, we emphasize that no existing approach combines dimension reduction via incremental basis updates with direct learning of reduced polynomial operators from streaming data, which is the contribution of our Streaming OpInf framework.
@@ 2.1 The Operator Inference Problem
    where the positive definite matrix $\boldsymbol{\Gamma}=\boldsymbol{\Gamma}^{\top}\in\mathbb{R}^{d\times d}$ with $d=r+r^{2}+m+1$ defines a Tikhonov regularization term McQuarrie et al. [2021].
@@
If $\mathbf{D}$ has full column rank and $\boldsymbol{\Gamma}\neq\mathbf{0}_{d\times d}$, this overdetermined system has unique solution $\mathbf{O}=\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{R}}\mskip-1.5mu}\mskip 1.5mu$. We present formulation (7) since it makes clear the RLS formulations introduced in Section 2.3.
Standard OpInf is a batch method built on standard direct computation of the SVD and direct solution of the LS problem. Loading and storing the entire data set in working memory incurs a memory cost of $O(n(K+r))$ for the SVD and $O(d(K+d))$ for the LS solution, which is infeasible for real-time streaming data settings and large-scale problems. While distributed computing Farcaş et al. [2025] and domain decomposition Farcas et al. [2024] address the memory bottleneck, we pursue an alternative by adapting OpInf to the streaming setting, where snapshots $\mathbf{x}_{1},\ldots,\mathbf{x}_{k}$ are processed incrementally as they arrive. This requires two key ingredients: (i) incremental basis construction via iSVD, and (ii) recursive operator learning via RLS which we now introduce.
@@ 2.2 Streaming SVD Methods
We seek to address the high memory cost in standard SVD for storing the data and basis, by introducing two streaming SVD algorithms, Baker’s iSVD Baker et al. [2012] and SketchySVD Tropp et al. [2019a].
@@ 2.2.1 Baker’s iSVD Algorithm
In this work, we employ Baker’s iSVD framework Baker et al. [2012], which builds on these identities (8) and (9) as summarized in Algorithm 1. Baker’s method is selected specifically because, similar to Zha and Simon [1999], it addresses two key challenges in standard iSVD methods: determining whether new snapshots warrant basis updates (typically via residual norm $p$ thresholding Brand [2002]) and maintaining numerical stability against accumulated errors (often requiring ad-hoc reorthogonalization Oxberry et al. [2017], Brand [2002]).
@@ 2.2.2 Randomized SketchySVD Algorithm
Let the sketch sizes be denoted $q$ and $s$, which control the dimensions of the compressed representations and determine the accuracy-memory tradeoff. SketchySVD maintains three randomized sketches $\mathcal{X}_{\mathsf{range}}\in\mathbb{R}^{n\times q},\,\mathcal{X}_{\mathsf{corange}}\in\mathbb{R}^{q\times K},\,\mathcal{X}_{\mathsf{core}}\in\mathbb{R}^{s\times s}$ that capture the range, corange, and core information of the data matrix $\mathbf{X}$, respectively Tropp et al. [2019a]. The core information is essential to improve the estimate of the singular values and vectors. These sketches are formed as products of $\mathbf{X}$ with four fixed random reduction maps $\boldsymbol{\Upsilon}\in\mathbb{R}^{q\times n},\,\boldsymbol{\Omega}\in\mathbb{R}^{q\times K},\,\boldsymbol{\Xi}\in\mathbb{R}^{s\times n},\,\boldsymbol{\Psi}\in\mathbb{R}^{s\times K}$:
@@
As summarized in Algorithm 2, SketchySVD processes each incoming snapshot $\mathbf{x}_{k}$ by updating the three sketches via efficient sparse matrix-vector multiplications, and then the random sketches are used to compute an estimate of the SVD only after all snapshots have been processed. To ensure a truncated SVD with sufficiently tight theoretical error bounds, the recommended sketch sizes are $q=4r+1$ and $s=2q+1$ Tropp et al. [2019a]. Further details are available in Tropp et al. [2019a].
@@ 2.2.3 Cost and Error Analysis
To assess suitability for our proposed method, we analyze memory and computational costs as well as sources of error for streaming SVD algorithms relative to batch SVD, tabulated in Table 1.
@@ 2.3 Recursive Least-Squares Methods
We introduce two RLS algorithms: standard RLS and a more numerically stable incremental QR-decomposition Recursive Least Squares (iQRRLS) in Sections 2.3.1 and 2.3.2, respectively.
@@ 2.3.2 Inverse QR-Decomposition RLS Algorithm
While the standard RLS algorithm 3 is effective for recursive parameter estimation, it suffers from numerical instabilities when updating the inverse correlation matrix $\mathbf{P}_{k}$ in finite-precision arithmetic. These instabilities arise because the update involves subtracting nearly equal positive definite matrices, causing catastrophic cancellation Sayed [2011]. QR-decomposition RLS algorithms address this by organizing iterated quantities and incoming data into arrays that are factorized using QR decomposition to update variables in a numerically stable manner.
These algorithms rely on transformations $\mathcal{A}_{k}\boldsymbol{\Theta}_{k}=\mathcal{B}_{k}$, where $\mathcal{A}_{k}$ is the pre-array containing previous quantities and new data, $\boldsymbol{\Theta}_{k}$ is a unitary matrix, and $\mathcal{B}_{k}$ is the triangular post-array containing updated quantities. This is essentially a QR decomposition of $\mathcal{A}_{k}^{\top}$. Two main variants exist: iQRRLS and QR-decomposition Recursive Least Squares (QRRLS), which differ in whether they update the inverse correlation matrix $\mathbf{P}_{k}$ or the correlation matrix $\sum_{i=1}^{k}\mathbf{d}_{i}^{\top}\mathbf{d}_{i}$ Sayed [2011], Haykin [2014].
The iQRRLS algorithm propagates the square root of the inverse correlation matrix $\mathbf{P}_{k}^{1/2}$ rather than $\mathbf{P}_{k}$ to ensure positive definiteness, using the Cholesky decomposition $\mathbf{P}_{k}=\mathbf{P}_{k}^{1/2}(\mathbf{P}_{k}^{1/2})^{\top\kern-0.5pt}$ where $\mathbf{P}_{k}^{1/2}$ is lower triangular. The algorithm organizes the transposed square root $\mathbf{P}_{k-1}^{\top/2}:=(\mathbf{P}_{k-1}^{1/2})^{\top\kern-0.5pt}$ and incoming data into the pre-array
@@
then performs QR decomposition to obtain the post-array from which updated quantities can be extracted. We summarize iQRRLS in Algorithm 4.
@@ 3.1 Streaming Reformulation of the Learning Step
To circumvent this additional data pass and its associated cost, we introduce a reformulation that expresses the LS data matrices directly in terms of the truncated SVD matrices obtained from iSVD, thereby avoiding explicit projection entirely. This reformulation approach yields the iSVD-LS and iSVD-RLS paradigms. While this reformulation eliminates the $O(nKr)$ projection cost and associated data pass, the projection-based paradigms offer their own advantages in terms of error control, as we elaborate in Section 3.3.1.
@@
$\begin{gathered}\mathbf{X}_{k}\odot\mathbf{X}_{k}\approx(\mathbf{V}_{r_{k}}\otimes\mathbf{V}_{r_{k}})(\boldsymbol{\Sigma}_{r_{k}}\otimes\boldsymbol{\Sigma}_{r_{k}})(\mathbf{W}_{r_{k}}^{\top\kern-0.5pt}\odot\mathbf{W}_{r_{k}}^{\top\kern-0.5pt})\implies\hat{\mathbf{X}}_{k}\odot\hat{\mathbf{X}}_{k}=\boldsymbol{\Sigma}_{r_{k}}\mathbf{W}_{r_{k}}^{\top}\odot\boldsymbol{\Sigma}_{r_{k}}\mathbf{W}_{r_{k}}^{\top},\end{gathered}$ (14)
where we use $(\mathbf{V}_{r_{k}}\otimes\mathbf{V}_{r_{k}})^{\top\kern-0.5pt}(\mathbf{V}_{r_{k}}\otimes\mathbf{V}_{r_{k}})=\mathbf{I}_{r_{k}^{2}}$.
Let $(i,j)$ denote the matrix entry at row $i$ and column $j$. Using (13) and (14), the data matrix $\mathbf{D}_{k}$ from (5) can be reformulated solely in terms of the rank-$r$ truncated matrices $\boldsymbol{\Sigma}_{r}=\boldsymbol{\Sigma}_{r_{k}}(1:r,1:r)$ and $\mathbf{W}_{r}=\mathbf{W}_{r_{k}}(:,1:r)$:
$\mathbf{D}_{k}=\begin{bmatrix}(\boldsymbol{\Sigma}_{r}\mathbf{W}_{r}^{\top})^{\top\kern-0.5pt}&(\boldsymbol{\Sigma}_{r}\mathbf{W}_{r}^{\top}\odot\boldsymbol{\Sigma}_{r}\mathbf{W}_{r}^{\top})^{\top\kern-0.5pt}&\mathbf{U}_{k}^{\top}&\mathbf{1}_{k}\end{bmatrix}$.
@@ 3.3.1 Errors from Streaming SVD
Baker’s iSVD produces deterministic errors that accumulate with each truncation Fareed and Singler [2020], Baker et al. [2012], while SketchySVD produces probabilistic errors controlled by sketch parameters Tropp et al. [2019a]. Both methods exhibit increased errors when spectral gaps diminish, as discussed in Section 2.2.3. These errors propagate to the learned operators through the LS data matrices. Specifically, the iSVD-LS/RLS paradigms construct data matrices (15) using right singular vectors $\mathbf{W}_{r}$ and singular values $\boldsymbol{\Sigma}_{r}$, which carry larger errors than left singular vectors according to (10). In contrast, the iSVD-Project-LS/RLS paradigms project data onto the POD basis using left singular vectors $\mathbf{V}_{r}$, yielding lower operator error at additional computational cost. We formalize this error propagation in the following lemma and theorem proven in Appendix C. Lemma 3.1 establishes a general perturbation bound for the learned operators in terms of perturbations to the data matrices.
@@
Let $\widetilde{\mathbf{D}}=\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu+[\mathbf{E}_{D}^{\top},\mathbf{0}_{d\times d}]^{\top\kern-0.5pt}$ and $\widetilde{\mathbf{R}}=\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{R}}\mskip-1.5mu}\mskip 1.5mu+[\mathbf{E}_{R}^{\top},\mathbf{0}_{r\times d}]^{\top\kern-0.5pt}$ denote perturbed data matrices with perturbations $\mathbf{E}_{D}\in\mathbb{R}^{K\times d}$ and $\mathbf{E}_{R}\in\mathbb{R}^{K\times r}$. Assume $\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu$ and $\widetilde{\mathbf{D}}$ both have full column rank $d$. Then the operator error between solutions $\mathbf{O}=\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{R}}\mskip-1.5mu}\mskip 1.5mu$ and $\widetilde{\mathbf{O}}=\widetilde{\mathbf{D}}^{\dagger}\widetilde{\mathbf{R}}$ satisfies
@@
We now combine Lemma 3.1 with explicit bounds on the data matrix perturbations $\mathbf{E}_{D}$ and $\mathbf{E}_{R}$ arising from streaming SVD errors to establish the following theorems, which bound the resulting operator errors for each paradigm.
Theorem 3.2 (Operator error bounds: Projection paradigms).
@@
Let $\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu$ and $\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{R}}\mskip-1.5mu}\mskip 1.5mu$ be constructed using the exact rank-$r$ SVD of $\mathbf{X}$, and let $\widetilde{\mathbf{D}}$ and $\widetilde{\mathbf{R}}$ be constructed using approximate factors $\widetilde{\mathbf{V}}_{r}$, $\widetilde{\boldsymbol{\Sigma}}_{r}$, $\widetilde{\mathbf{W}}_{r}$ from streaming SVD with the left singular vector error $\tau_{v}=\|\mathbf{V}_{r}-\widetilde{\mathbf{V}}_{r}\|_{F}$. Assume the time-derivative data is computed via finite differences, $\dot{\mathbf{X}}=\mathbf{X}\boldsymbol{\Delta}_{K}$. Furthermore, suppose $\|\mathbf{X}-\widetilde{\mathbf{X}}\|_{2}\leq\varepsilon$ and $\boldsymbol{\Gamma}=\mathrm{diag}(\gamma_{1},\ldots,\gamma_{d})$ with $\gamma_{j}>0$ for all $j$, and let $\sqrt{\gamma_{\min}}=\min_{j}\sqrt{\gamma_{j}}$. Then for the iSVD-Project-LS/RLS paradigms, the data matrix perturbations satisfy